OpenStudy (anikate):

find the common difference, d, for the arithmetic sequence in which a sub1=9 and a sub 20=-29

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anikate):

@Calcmathlete

OpenStudy (anonymous):

Use the formula: \[a_n = a_1 + (n - 1)d\]

OpenStudy (anikate):

how?

OpenStudy (anikate):

can you plz show what to substitute

OpenStudy (anonymous):

\[a_n = a_1 + (n - 1)d\]\[a_{20} = a_1 + (20 - 1)d\]\[-29 = 9 + 19d\]Solve for d.

OpenStudy (anikate):

-2

OpenStudy (anonymous):

Yup. d = -2

OpenStudy (anikate):

one more

OpenStudy (anikate):

you plz help me with one more?

OpenStudy (anonymous):

Alright

OpenStudy (anikate):

In geometric sequence \[a _{3}=2\] and \[r=1/2\] find \[a _{8}\]

OpenStudy (anonymous):

Do you know how to do this one?

OpenStudy (anikate):

nope not a bit

OpenStudy (anonymous):

Use the formula \[a_n = a_1r^{n - 1}\]

OpenStudy (anikate):

whta do i substitue for a1

OpenStudy (anonymous):

First solve for \(a_1\). \[a_3 = a_1(\frac12)^2\]\[2 = a_1(\frac12)^2\]Can you solve?

OpenStudy (anikate):

a1= 8?

OpenStudy (anonymous):

Yup...

OpenStudy (anikate):

y? 1/2 ^2

OpenStudy (anonymous):

That's the formula... \[a_n = a_1r^{n - 1}\]\[a_3 = a_1(r)^{3 - 1}\]\[2 = a_1(\frac12)^2\]

OpenStudy (anikate):

h ok

OpenStudy (anikate):

alright I cant take it from here thx calc

OpenStudy (anonymous):

Alright :)

OpenStudy (anikate):

may I have a a medal if its alright with you :)

OpenStudy (anikate):

thx :D

OpenStudy (anonymous):

Wait, depends on what your answer is XD

OpenStudy (anikate):

I'll you tom. right now I gt study

OpenStudy (anonymous):

lol k :)